OpenStudy (anonymous):

In how many ways can 8 people each have different birthdays? Assume that there are 365 days in a year.

OpenStudy (anonymous):

365*364*363*362*361*360*359*358 or \[\frac{365!}{358!}\]

OpenStudy (anonymous):

how did you know how to do that?

OpenStudy (anonymous):

are you sure 365 -8 =357 ?

OpenStudy (anonymous):

sorry should have been \[\frac{365!}{357!}\]

OpenStudy (anonymous):

why 357?

OpenStudy (anonymous):

first person can have birthday any day of the year, 365 days second person , every day except ^ that one day 364 ................ ............... .................

OpenStudy (anonymous):

gotchaaaa